Seminar on mirror symmetry and representation theory

(upgraded from “Seminar on derived categories in Algebraic Geometry”)

Tuesdays, 5.00-6.30, MIT, room 2-132.

Tuesday, February 21

Homological mirror symmetry for ℂℙ^{2}.

by Denis Auroux (MIT).

The goal of this talk will be to explain the statement of Kontsevich’s homological mirror symmetry conjecture in the case of Fano varieties. We will in particular attempt to give a rough definition of the derived category of Lagrangian vanishing cycles of a Landau-Ginzburg model. ℂℙ^{2} will be used as our main example to illustrate the various derived

categories which enter the picture.

seminar contacts: Roman Bezrukavnikov (bezrukav@math.mit.edu); Rina Anno (rina@math.harvard.edu)